40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1361     Entry time: Tue Oct 8 23:01:56 2013
Author: Evan 
Type: DailyProgress 
Category: BEAT 
Subject: Attempts at new beat measurement 

[Tara, Evan]

Having successfully floated the table yesterday, we attempted a new beat measurement in the hopes that the large shelf below 100 Hz had disappeared. Unfortunately, this appears to not be the case. Additionally, many of our signals are plagued by unusually large, slow drifts. We're hoping that they're just thermal transients caused by all the work on the table over the past 12 hours, and that by tomorrow things will have settled down. We'll see if that's the case.

Anyway, we did the following things today:

  • We reconnected cables that come in from off the table and go onto cameras, PDs, etc., paying special attention to strain relief and vibration isolation since the table now floats.
  • We redid the alignment to recover ~90% visibility; this required only touching the periscope mirrors (somewhat surprising considering what we subjected the table to in order to switch out the legs).
  • We got the cavity PDH loops up and running again. The control signals show unusually large drifts. We also noticed this while sweeping the lasers to align the cavities; the resonance for north in particular would wander out of the sweep range every 30 seconds even though the laser was being driven at 10 Vpp from an SRS function generator.
  • We spent some time trying to null (what we assume is) RFAM-induced offset in the PDH error signals. We did this by adjusting the HWPs before each cavity EOM and nulling the offset on TTFSS common OUT1. The south cavity already had a small offset, so no adjustment was required. On the north cavity, there was a noticeable offset (~20 mV, compared to an error signal pk-pk of 220 mV), so Tara nulled it. We then found that we could get a stable lock with the laser PZT actuator alone, and that adding the EOM actuator caused the loop to oscillate (almost as if the EOM actuator was driven with the wrong sign). So we looked at the error signal again, and unexpectedly found another ~20 mV offset. Tara nulled it again and this time the lock was stable; in fact, we were able to get the common and fast gain knobs up to 1000 and 1000 (compared to 800 and 800 earlier in the day). No idea what the problem is here; possibly it drifted between the successive adjustments.
  • We looked at the beat. It appears to be not much better than with the table unfloated.
  • We took a measurement of RIN-induced disturbance in the beat by driving the south EOAM with a sine from the SR785 and taking the TF that takes transmission PD intensity to beat fluctuation. Unfortunately, this measurement is not consistent (in magnitude or phase) between successive sweeps. It seems to be due (at least in part) to DC drift in the beat.
  • We tried turning on the crude south ISS, but it made the beat more noisy.

 

ELOG V3.1.3-